# Occlusion-aware Text-Image-Point Cloud Pretraining for Open-World 3D Object Recognition

### v1

@@ Abstract
Recent open-world representation learning approaches have leveraged CLIP to enable zero-shot 3D object recognition. However, performance on real point clouds with occlusions still falls short due to the unrealistic pretraining settings. Additionally, these methods incur high inference costs because they rely on Transformer’s attention modules. In this paper, we make two contributions to address these limitations. First, we propose occlusion-aware text-image-point cloud pretraining to reduce the training-testing domain gap. From 52K synthetic 3D objects, our framework generates nearly 630K partial point clouds for pretraining, consistently improving real-world recognition performances of existing popular 3D networks. Second, to reduce computational requirements, we introduce DuoMamba, a two-stream linear state space model tailored for point clouds. By integrating two space-filling curves with 1D convolutions, DuoMamba effectively models spatial dependencies between point tokens, offering a powerful alternative to Transformer. When pretrained with our framework, DuoMamba surpasses current state-of-the-art methods while reducing latency and FLOPs, highlighting the potential of our approach for real-world applications. We will release our data and code to facilitate future research.
@@ 1 Introduction
Existing works in this line of research take 3D-image-text triplets as input and align the three embedding spaces using cross-modal contrastive learning. These methods represent 3D shapes either as depth maps [26, 16, 53, 59] or raw point clouds [23, 54, 48, 49, 57, 10]. Depth-based approaches must first convert point clouds into 2D depth maps and use pretrained image encoders, such as Vision Transformer (ViT) [9], for 3D feature extractions. However, their performance typically suffers from information loss during the projection and the domain gap caused by differences between RGB and depth images. On the other hand, point-based methods [48, 23, 49, 57, 10, 54] can directly exploit all intrinsic geometry in the point clouds. An example is the recent work CLIP2 [52], which pretrains a 3D encoder using real-scanned objects extracted from scene-level point clouds. For contrastive learning, it pairs these with cropped images and simple category-based prompts as text descriptions. However, due to limited diversity in captions and the low quality of cropped images (caused by occlusion, lighting condition, image resolution, etc), this method faces challenges in transferring CLIP’s knowledge, resulting in suboptimal performance.
@@
Other works [48, 49, 23, 57, 10, 54] instead leverage synthetic 3D models [^1] In this paper, we use 3D models to refer to 3D CAD models or 3D meshes instead of 3D deep learning models. to construct pretraining triplets. These methods uniformly sample points from the mesh surface to create full point clouds [^2] A full (complete) point cloud provides 360-degree coverage of an object, while a partial (occluded) one is captured from a single viewpoint.. They also render RGB images from preset camera positions and generate diverse captions from multiple sources, allowing for control over the quality of images and texts. As a result, these methods demonstrate promising zero-shot performance on complete point cloud benchmarks such as ModelNet40 [47]. However, their performance degrades significantly on real-scanned data, leading to unsatisfactory results in practical scenarios. As shown in Figure 1a, there is a 20% accuracy drop from the synthetic ModelNet40 [47] to the real ScanObjectNN [41], caused by the large domain gap between complete point clouds in pretraining and occluded ones encountered in real-world conditions. To address this data discrepancy, we introduce an occlusion-aware pretraining framework that leverages synthetic 3D meshes to create partial point clouds. We simulate real-world scenarios by putting a virtual camera around an object and only sample points visible from the camera position. From 52K ShapeNetCore [2] 3D models, our framework generates nearly 630K occluded point clouds for pretraining and enhances the zero-shot accuracy of SparseConv [4] and PointBERT [50] by 3.5% and 5.1% on ScanObjectNN [41]. Despite using only synthetic objects, our framework consistently improves recognition performance on various real-world tasks and even outperforms methods that use real-scanned data.
@@ 6.4 Zero-Shot 3D Object Detection
(c) Mamba-based encoders comparison. Our DuoMamba achieves a better computation-performance trade-off than previous methods.
@@ Acknowledgment.
This research was supported by the Australian Research Council (ARC) under the discovery grant project #240101926. Khanh Nguyen is supported by the tuition fee and stipend scholarships from the University of Western Australia. Professor Ajmal Mian is the recipient of an ARC Future Fellowship Award (project #FT210100268) funded by the Australian Government.

### v2

@@ Abstract
Recent open-world representation learning approaches have leveraged CLIP to enable zero-shot 3D object recognition. However, performance on real point clouds with occlusions still falls short due to unrealistic pretraining settings. Additionally, these methods incur high inference costs because they rely on Transformer’s attention modules. In this paper, we make two contributions to address these limitations. First, we propose occlusion-aware text-image-point cloud pretraining to reduce the training-testing domain gap. From 52K synthetic 3D objects, our framework generates nearly 630K partial point clouds for pretraining, consistently improving real-world recognition performances of existing popular 3D networks. Second, to reduce computational requirements, we introduce DuoMamba, a two-stream linear state space model tailored for point clouds. By integrating two space-filling curves with 1D convolutions, DuoMamba effectively models spatial dependencies between point tokens, offering a powerful alternative to Transformer. When pretrained with our framework, DuoMamba surpasses current state-of-the-art methods while reducing latency and FLOPs, highlighting the potential of our approach for real-world applications. Our code and data are available at ndkhanh360.github.io/project-occtip.
@@ 1 Introduction
Existing works in this line of research take 3D-image-text triplets as input and align the three embedding spaces using cross-modal contrastive learning. These methods represent 3D shapes either as depth maps [29, 19, 58, 65] or raw point clouds [26, 59, 53, 54, 62, 12]. Depth-based approaches must first convert point clouds into 2D depth maps and use pretrained image encoders, such as Vision Transformer (ViT) [10], for 3D feature extractions. However, their performance typically suffers from information loss during the projection and the domain gap caused by differences between RGB and depth images. On the other hand, point-based methods [53, 26, 54, 62, 12, 59] can directly exploit all intrinsic geometry in the point clouds. An example is the recent work CLIP2 [57], which pretrains a 3D encoder using real-scanned objects extracted from scene-level point clouds. For contrastive learning, it pairs these with cropped images and simple category-based prompts as text descriptions. However, limited caption diversity and poor cropped image quality (due to occlusion, lighting, etc.) hinder CLIP knowledge transfer, leading to suboptimal performance.
@@
Other works [53, 54, 26, 62, 12, 59] instead leverage synthetic 3D models [^1] In this paper, we use 3D models to refer to 3D CAD models or 3D meshes instead of 3D deep learning models. to construct pretraining triplets. These methods uniformly sample points from the mesh surface to create full point clouds [^2] A full (complete) point cloud provides 360-degree coverage of an object, while a partial (occluded) one is captured from a single viewpoint.. They also render RGB images from preset camera positions and generate diverse captions from multiple sources, allowing for control over the quality of images and texts. As a result, these methods demonstrate promising zero-shot performance on complete point cloud benchmarks such as ModelNet40 [52]. However, their performance degrades significantly on real-scanned data, leading to unsatisfactory results in practical scenarios. As shown in Figure 1a, there is a 20% accuracy drop from the synthetic ModelNet40 [52] to the real ScanObjectNN [46], caused by the large domain gap between complete point clouds in pretraining and occluded ones encountered in real-world conditions. To address this data discrepancy, we introduce an occlusion-aware pretraining framework that leverages synthetic 3D meshes to create partial point clouds. We simulate real-world scenarios by putting a virtual camera around an object and only sample points visible from the camera position. From 52K ShapeNetCore [2] 3D models, our framework generates nearly 630K occluded point clouds for pretraining and enhances the zero-shot accuracy of SparseConv [4] and PointBERT [55] by 3.8% and 5.1% on ScanObjectNN [46]. Despite using only synthetic objects, our framework consistently improves recognition performance on various real-world tasks and even outperforms methods that use real-scanned data.
@@ 6.4 Zero-Shot 3D Object Detection
(c) Mamba-based encoders comparison. Our DuoMamba architecture achieves a better computation-performance trade-off than previous methods.
@@ Acknowledgment.
The authors sincerely thank Mr. Trong-Thang Pham for his valuable feedback during the preparation of this paper. This research was supported by the Australian Research Council (ARC) under discovery grant project #240101926. Professor Ajmal Mian is the recipient of an ARC Future Fellowship Award (project #FT210100268) funded by the Australian Government. Mr. Khanh Nguyen is supported by the tuition fee scholarship from the University of Western Australia and a stipend from the ARC project #FT210100268.
Supplementary Material
8 Additional Discussion on Existing Works
Discussion on Occlusion Methods.
Ren et al. [41] simplified occlusion by treating it as a form of corruption, referred to as “Drop Local,” where k-NN clusters are randomly removed from point clouds. They then proposed an architecture and an augmentation strategy (based on deforming and mixing objects) to address general corruptions rather than focusing on occlusion. Hamdi et al. [16] introduced a viewpoint prediction module as a component for multi-view 3D recognition (which rely on 3D-to-2D projection). By predicting ‘good’ views to render images from point clouds, indirectly, the recognition model becomes more robust to occlusion (empirically simulated by randomly cropping the object point clouds along canonical directions). In contrast, our OccTIP method more realistically simulates self-occlusion through the rendering process and integrates single-view point clouds during pretraining, improving occlusion robustness for any point cloud encoders.
Comparison with VisionMamba (Vim).
While Vim [64] also has a two-stream design, it has two key limitations: (1) reliance on one-directional neighborhood aggregation (CausalConv1D) and (2) only able to utilize a single neighborhood structure due to its simple forward and backward scanning strategy. In contrast, DuoMamba uses Conv1D for bidirectional local aggregation and can flexibly process two diverse orderings (e.g., Hilbert, Trans-Hilbert) simultaneously within a single block to fully exploit 3D geometry of the point clouds. These technical enhancements lead to improved performance as shown in Table 6.
Dataset Vim [64] Vim [64] + Hilbert DuoMamba
ModelNet40-P 65.3 63.8 67.7
ScanObjectNN 61.1 62.7 63.5
Table 6: Zero-shot accuracy of Vim and DuoMamba.
9 Implementation Details
Triplet Generations.
We render RBG images with a resolution of $512\times 512$ and a transparent background. Similar to OpenShape [26], descriptions for each object come from three sources: (1) raw texts from the dataset’s metadata, (2) captions generated by BLIP [24] and Azure Cognitive Services, (3) retrieved captions from visually similar images in the LAION-5B [43] dataset. The first source of captions (created from metadata) includes three texts: (a) object name, (b) object category, and (c) concatenation of the subcategory name.
Training Details.
During pretraining, we use a batch size of 32 and randomly replace point colors with a constant value of 0.4 with a probability of 0.5. During testing, we assign the same constant value to point clouds that do not have color information, such as those in the ScanObjectNN [46] dataset. For more efficient training, we precompute and cache text and image features from CLIP [39] and directly use them as inputs to the text and image projection heads. Since there is significant fluctuation when training with partial point clouds, we follow [12] to employ Exponential Moving Average (EMA) [45] with a decay factor of 0.9995 to stabilize the training process. We use a cosine learning rate scheduler with a base learning rate of $7\mathrm{e}{-4}$.
Method Mean Cab Bed Chair Sofa Tabl Door Wind Bksf Pic Cntr Desk Curt Fridg ShwrCurt Toil Sink Bath Bin
AP25 PointCLIP [19] 6.00 3.99 4.82 45.16 4.82 7.36 4.62 2.19 - - 1.02 4.00 - - - - 13.40 6.46 -
PointCLIP V2 [65] 18.97 19.32 20.98 61.89 15.55 23.78 13.22 17.42 - - 12.43 21.43 - - - - 14.54 16.77 -
OpenShape [26] 20.40 9.63 38.62 73.05 57.28 37.00 29.52 5.74 23.94 2.07 3.37 16.25 1.25 4.45 0.84 9.00 22.76 16.21 16.23
MixCon3D [12] 24.11 11.55 43.21 79.33 63.97 42.91 29.94 4.85 25.26 3.98 1.49 25.58 2.00 4.95 0.81 13.23 20.58 38.03 22.25
TAMM [59] 23.07 10.03 32.68 75.16 55.73 36.72 32.44 5.26 24.82 2.52 2.04 22.53 2.11 3.26 1.23 17.83 23.87 46.50 20.48
OccTIP 28.92 12.85 56.43 80.41 68.78 40.11 37.68 7.09 30.51 3.21 2.46 31.55 5.18 8.54 2.14 29.89 35.64 41.93 26.24
AP50 PointCLIP [58] 4.76 1.67 4.33 39.53 3.65 5.97 2.61 0.52 - - 0.42 2.45 - - - - 5.27 1.31 -
PointCLIP V2 [65] 11.53 10.43 13.54 41.23 6.60 15.21 6.23 11.35 - - 6.23 10.84 - - - - 11.43 10.14 -
OpenShape [26] 16.12 3.78 36.99 62.48 49.48 33.05 17.40 2.12 21.97 0.61 1.34 11.97 0.45 4.18 0.59 8.38 10.68 16.16 8.55
MixCon3D [12] 19.09 3.61 41.90 67.67 51.13 38.22 17.34 1.56 23.44 1.56 0.36 18.63 0.59 4.71 0.43 12.07 9.18 37.69 13.51
TAMM [59] 18.11 3.10 31.64 64.35 42.51 30.82 20.55 2.11 21.26 0.85 0.50 17.71 0.80 3.09 0.81 17.00 10.44 46.27 12.26
OccTIP 22.73 5.44 54.77 68.91 55.53 34.55 22.55 2.92 25.71 0.98 0.84 22.91 2.34 8.36 1.31 27.27 16.86 41.65 16.27
Table 7: Zero-shot 3D object detection results on ScanNetV2 [6]. Our method OccTIP achieves the highest mAP and consistently has the highest or second-highest AP scores across most categories, showing the superiority of the proposed approach in complex real-world recognition. (*: results obtained using released pretrained weights, : results reproduced using the authors’ public code.)
Method Mean Bed Table Sofa Chair Toilet Desk Dresser Night Stand Bookshelf Bathtub
AP25 OpenShape [26] 18.61 33.09 24.18 28.96 45.51 10.42 13.58 2.75 11.77 11.13 4.71
MixCon3D [12] 18.69 28.25 26.75 34.44 47.77 6.05 15.76 2.31 11.56 6.91 7.14
TAMM [59] 18.91 18.15 27.78 27.67 47.00 21.41 14.54 2.43 10.81 11.14 8.20
OccTIP 24.37 43.45 29.21 34.22 51.19 12.78 18.16 3.76 11.14 13.96 25.90
AP50 OpenShape [26] 9.78 23.71 9.01 20.85 24.37 7.74 3.02 1.00 5.47 1.77 0.89
MixCon3D [12] 9.63 17.97 10.22 24.53 26.00 3.80 3.38 0.51 6.30 1.73 1.86
TAMM [59] 9.96 12.37 11.01 20.36 25.41 17.96 3.22 0.81 4.87 1.71 1.90
OccTIP 13.01 32.67 11.21 25.46 28.04 8.50 4.33 1.71 5.11 1.92 11.18
Table 8: Zero-shot 3D object detection results on SUN RGB-D [44]. Our method OccTIP achieves the highest mAP and consistently has the highest or second-highest AP scores across most categories, showing the superiority of the proposed approach in complex real-world recognition. (*: results obtained using released pretrained weights, : results reproduced using the authors’ public code.)
10 Comparisons with Previous Works Pretrained on Larger Datasets
We further compare our method (pretrained on 52K ShapeNetCore [2] objects) with previous works pretrained on a significantly larger ensemble of 880K 3D objects from four datasets: ShapeNetCore [2], ABO [5], 3D-FUTURE [11], and Objaverse [7]. We use the official results reported in previous papers and evaluate all approaches on the real-world ScanObjectNN [46] dataset to assess their recognition performance in practical scenarios.
Model Size and Zero-Shot Object Classification Performance.
We compare the parameter counts of various point cloud encoders and their zero-shot performance in Figure 5. Despite only being pretrained on ShapeNetCore [2], our DuoMamba outperforms all existing models of comparable size that are pretrained on 880K 3D objects – 17 times more data. Notably, the zero-shot accuracy gap between our model and the best-performing model Uni3D-giant [62] is just 1.8%, even though our model is only 1/35 its size. This highlights DuoMamba’s superior size-to-performance efficiency. Scaling up the model and pretraining on larger datasets is likely to further enhance performance, which we leave as future work.
Few-Shot Linear Probing.
We perform a few-shot experiment similar to the one in Section 6.2 (main paper), this time comparing our approach against models pretrained on the ensemble of 880K 3D objects. As illustrated in Figure 6, our method consistently outperforms all other works across all few-shot settings, highlighting our pretraining framework’s data efficiency and effectiveness in learning robust and generalizable features for real-world recognition.
Figure 5: Comparisons of model size and zero-shot accuracy on ScanObjectNN [46]. Our model is pretrained on 52K ShapeNetCore [2] objects, whereas all other approaches are pretrained on an ensemble of 880K objects from four datasets: Objaverse [7], ABO [5], 3D-FUTURE [11], and ShapeNetCore [2]. Despite being pretrained on a less diverse set of objects and having the smallest size, DuoMamba demonstrates competitive performance. Among models with fewer than 50M parameters (DuoMamba, PointBERT [55], SparseConv [4]), our model outperforms all others by a significant margin of 3% in zero-shot accuracy. While Uni3D-giant [62] achieves a slightly higher accuracy with a gap of 1.8%, it comes at the cost of a substantially larger model size, with 1016.5M parameters – 35 times the size of DuoMamba. This highlights the optimal balance between model size and performance offered by our method compared to existing approaches.
Figure 6: Few-shot linear probing on ScanObjectNN [46]. Our method is pretrained on 52K ShapeNetCore [2] objects, whereas other models are pretrained on 880K objects. Despite using significantly less data, our framework OccTIP outperforms all existing methods across all few-shot settings, demonstrating the data efficiency and the high-quality latent space learned by our approach.
11 Additional Quantitative Results
Evaluate Pretrained DuoMamba on ModelNet40.
To evaluate DuoMamba (pretrained with OccTIP) on complete point clouds, we generate partial point clouds from 12 views (as in pretraining) and use majority voting for class prediction. Figure 7 shows that on ModelNet40, we perform competitively with previous works pretrained on full point clouds and even surpass OpenShape by 1.3%.
Figure 7: Comparison with methods pretrained on complete point clouds.
Complete Results for Zero-Shot 3D Object Detection.
The average precision (AP) for each class and the mean Average Precision (mAP) for the zero-shot 3D object detection experiments (Section 6.4 in the main paper) are provided in Table 7 (for ScanNetV2 [6] benchmark) and Table 8 (for SUN RGB-D [44] benchmark). Our method OccTIP consistently achieves the best or second-best AP across most categories and achieves the highest mAP, with a significant margin over existing techniques on both datasets. These results highlight the effectiveness of OccTIP and its applicability to complex, real-world recognition tasks.
Pretraining with Complete vs. Partial Point Clouds.
Table 9 shows that our synthetic partial data consistently improves all models’ accuracy on real-world ScanObjectNN, with DuoMamba performing best in both settings.
Pretraining data SparseConv PointBERT DuoMamba
Complete 56.0 55.5 57.5
Partial (OccTIP) 61.7 (+5.7) 60.6 (+5.1) 63.5 (+6.0)
Table 9: ScanObjectNN accuracy when pretraining with full vs partial data.
Architecture Influence on Object Detection Performance.
Table 10 compares object detection performance of DuoMamba and PointBERT pretrained with OccTIP against PointBERT’s best performance by previous pretraining baselines. OccTIP consistently enhances PointBERT’s performance, and its combination with DuoMamba achieves the best results.
Pretraining Encoder ScanNetV2 SUN RGB-D
method mAP25 mAP50 mAP25 mAP50
Best current PointBERT 24.1 19.1 18.9 10.0
OccTIP 25.4 19.3 21.9 11.7
OccTIP DuoMamba 28.9 22.7 24.4 13.0
Table 10: Detection results of different models and pretraining methods.